OpenStudy (anonymous):

Simplify 3^2 4 / 3 7 9.

OpenStudy (anonymous):

is it \[\frac{3^2\times 4}{3\times 7\times 9}\]??

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

ok then since \(3^2=9\) you have \[\frac{3^2\times 4}{3\times 7\times 9}=\frac{9\times 4}{3\times 7\times 9}\] \[=\frac{\cancel{9}\times 4}{3\times 7\times \cancel{9}}=\frac{4}{3\times 7}\]

OpenStudy (anonymous):

your last job is to multiply \(3\times 7\) in the denominator and you are done

OpenStudy (anonymous):

Thank you so much!(:

OpenStudy (anonymous):

yw